OpenStudy (realitywillslapyou):

Which equation is not equivalent to the formula d = rt? A. r = (t/d) B. r = (d/t) C. t= (d/r) D. d = tr

OpenStudy (anonymous):

ouch u slapped me

OpenStudy (realitywillslapyou):

When did I do that?

OpenStudy (anonymous):

right now

OpenStudy (realitywillslapyou):

m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m m <---- my cat typed that

OpenStudy (anonymous):

i can tell

OpenStudy (realitywillslapyou):

lol XD

OpenStudy (anonymous):

but still that hurt

OpenStudy (realitywillslapyou):

I'm sorry sometimes I just randomly slap things and it just so happened that I slapped you. :P

OpenStudy (anonymous):

=D

OpenStudy (realitywillslapyou):

So, are you going to help me with my question?

OpenStudy (anonymous):

ill try

OpenStudy (realitywillslapyou):

Okay, thank you :)

OpenStudy (anonymous):

the answer is A. r = (t/d)

OpenStudy (realitywillslapyou):

Yay! Can you please help me with some more?

OpenStudy (realitywillslapyou):

I promise I won't slap you!

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

but u already slaped me

OpenStudy (realitywillslapyou):

I won't slap you again, and I explained it to you I just slap randomly! Consider the functions f(x) = 2(x − 3) and g(x) = 3x. What is the value of (f + g)(3)? A. 0 B. 9 C. −9 D. 2

OpenStudy (realitywillslapyou):

Blame bad luck! :P

OpenStudy (anonymous):

=D

OpenStudy (anonymous):

sorry i dont know =(

OpenStudy (realitywillslapyou):

That's okay.....

OpenStudy (realitywillslapyou):

Campbell Music Academy held workshops on classical music composers Mozart, Vivaldi, and Haydn. The Venn diagram shows the number of seniors who chose to attend workshops on each of the three composers. If the total number of seniors in the academy is 64, what is the total number of students who belong to the set V ∩ (~H ∩ ~ M)? A. 45 B. 33 C. 19 D. 15

OpenStudy (anonymous):

first one is ~H second one is ~M third one is what you are looking for in english, "in V and not in H and not in M" so 15

OpenStudy (realitywillslapyou):

Thankies! :D

OpenStudy (anonymous):

ur welcome

OpenStudy (realitywillslapyou):

What is the equation of the line that passes through the point (3, 2) and has a slope of 4? A. y = 4x − 1 B. y = 4x − 5 C. y = 4x − 10 D. y = 4x + 1

OpenStudy (anonymous):

@bob-bob-bob quit spamming and actualy help people

OpenStudy (anonymous):

i did now MIND UR OWN BUSINESS